OpenStudy (anonymous):

. Robert has 12 red folders, 5 blue folders, 13 green folders, and 9 black folders on his shelf. If he randomly grabs a folder, what is the probability that it will be either green or black?

OpenStudy (swissgirl):

He has a total of (12+5+13+9)=39 Folder The probability that Robert will choose a folder that is either red or green is The probability that he will choose a green folder plus the probability that he will choose a red folder. The probability that he will choose a red folder is the amount of red folders divided by the total amount of folders and the same is with green folders Therefore we have the following equation Probability that will choose red or green folder =(13/39)+(12/39)=(13+12)/39=25/39